OpenStudy (anonymous):

find the value for x x-8y=20 x+11y=-21

OpenStudy (diyadiya):

x-8y=20 x+11y=-21 Subtratct both the equations What do you get?

OpenStudy (anonymous):

so x is -3

OpenStudy (ajprincess):

x-8y=20-(1) x+11y=-21-(2) (2)-(1)=>19y=-41 y=-41/19 (1)=>x=20+8y =20-(8*41/19) =(20*19-328)/19 =(380-328)/19 =52/19

OpenStudy (anonymous):

so what is x

OpenStudy (ajprincess):

x=52/19 =2.73684 Which will be 3 when rounded off

OpenStudy (anonymous):

so what is y

OpenStudy (diyadiya):

@ajprincess solved it completely .. what's x & y is very clear

OpenStudy (ajprincess):

y=-41/19 =2.15789 which is equal to 2 when rounded off